# Twinkle Twinkle Little Star, Roman Sees Where You Are: Predicting Exoplanet Transit Yields in the Rosette Nebula with the Nancy Grace Roman Space Telescope

### v1

@@ 6 Summary
Our simulations incorporate spectral-type-dependent occurrence rates, age-dependent radius inflation, and a conservative $7\sigma$ detection threshold. We predict the detection of $33\pm 9$ transiting exoplanets in a month-long survey and $29\pm 8$ exoplanets in a two-week survey, with approximately 90% orbiting M dwarfs. The extended baseline primarily improves sensitivity to longer-period planets orbiting FGK stars, while most M dwarf detections are captured within two weeks. These detections are dominated by of 1-2 $R_{\oplus}$ super-Earths and sub-Neptunes with periods less than 8 days, representing an order-of-magnitude increase over the current sample of only three confirmed planets younger than 20 Myr.
@@
We gratefully acknowledge the generous support of NASA under Award No. UGS25_12-1 through the Wisconsin Space Grant Consortium. We also express gratitude to support from the Peter Livingston Scholars Program, whose contributions to undergraduate research have played a key role in the development of this work. Support for this research was provided by the Office of the Vice Chancellor for Research and Graduate Education at the University of Wisconsin–Madison with funding from the Wisconsin Alumni Research Foundation. For our computational needs, our pipeline used HTCondor, a high-throughput computing cluster offered by the Center for High Throughput Computing (CHTC) at the University of Wisconsin-Madison. R.S.N. gratefully acknowledges Adam Distler for insightful discussions and comments during the preliminary development of this work’s methodologies and analysis framework. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work has also made use of the NASA Exoplanet Archive (http://dx.doi.org/10.26133/NEA1 (catalog 10.26133/NEA1)), which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This work was conducted at the University of Wisconsin-Madison, which is located on occupied ancestral land of the Ho-Chunk people, a place their nation has called Teejop since time immemorial. In an 1832 treaty, the Ho-Chunk were forced to cede this territory. The university was founded on and funded through this seized land; this legacy enabled the science presented here.

### v2

@@ VI Summary
Our simulations incorporate spectral-type-dependent occurrence rates, age-dependent radius inflation, and a conservative $7\sigma$ detection threshold. We predict the detection of $33\pm 9$ transiting exoplanets in a month-long survey and $29\pm 8$ exoplanets in a two-week survey, with approximately 90% orbiting M dwarfs. The extended baseline primarily improves sensitivity to longer-period planets orbiting FGK stars, while most M dwarf detections are captured within two weeks. These detections are dominated by of 1-2 $R_{\oplus}$ super-Earths and sub-Neptunes with periods less than 8 days, representing an order-of-magnitude increase over the current sample of only three confirmed transiting planets younger than 20 Myr around FGKM stars.
@@
We are grateful to the anonymous reviewer for providing thoughtful comments that improved the scientific analyses presented in this paper. We gratefully acknowledge the generous support of NASA under Award No. UGS25_12-1 through the Wisconsin Space Grant Consortium. We also express gratitude to support from the Peter Livingston Scholars Program, whose contributions to undergraduate research have played a key role in the development of this work. Support for this research was provided by the Office of the Vice Chancellor for Research and Graduate Education at the University of Wisconsin–Madison with funding from the Wisconsin Alumni Research Foundation. For our computational needs, our pipeline used HTCondor, a high-throughput computing cluster offered by the Center for High Throughput Computing (CHTC) at the University of Wisconsin-Madison. R.S.N. gratefully acknowledges Adam Distler for insightful discussions and comments during the preliminary development of this work’s methodologies and analysis framework. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work has also made use of the NASA Exoplanet Archive (http://dx.doi.org/10.26133/NEA1 (catalog 10.26133/NEA1)), which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This work was conducted at the University of Wisconsin-Madison, which is located on occupied ancestral land of the Ho-Chunk people, a place their nation has called Teejop since time immemorial. In an 1832 treaty, the Ho-Chunk were forced to cede this territory. The university was founded on and funded through this seized land; this legacy enabled the science presented here. JMV acknowledges support from the European Union through the Exo-PEA ERC project (grant number 101164652). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
We present normalized Kernel Density Estimates for the observed distributions of detected planets in our simulations.
Figure 9: Normalized Kernel Density Estimates for detected planet candidates orbiting FGK stars (left) and M stars (right) comparing one-month (solid) and two-week (dashed) survey baselines. Rows show the distributions of orbital period, planet radius, and host stellar mass (with corresponding dust-corrected F146 magnitude on the secondary axis).